# Task Graph offloading via Deep Reinforcement Learning in Mobile Edge Computing

### v1

@@ 6.3 Evaluation for Reinforcement Learning
In this section, we use cumulative rewards to evaluate the performance of our reinforcement learning. Considering that MUs can randomly send offloading requests of applications, we evaluate cumulative rewards of the reinforcement learning with different arrival rates, i.e., $\lambda=\{5,7,9\}$, when it handles 10 applications. In the experiments, we let 10 applications per episode reach the simulator with different arrival rates, then launch Algorithm 3 to train the agent and observe the reward values. From Figs. 5, we can observe that the total rewards obtained by the agent are not large at the beginning of the learning process. The total rewards increase as we augment the number of episodes for training. Moreover, the total rewards show relatively stable trends when agent is trained for more than 600 episodes, which is shown in Fig. 5(c). This can be interpreted that the agent can make a relatively optimized decision in every scheduling for the task. As a result, the cumulative rewards can converge. The above results demonstrate the convergence performance of SATA-DRL.
@@ 6.4 Evaluation for Optimization Objective
To evaluate the optimization performance of SATA-DRL, we compare it with HEFTtopcuoglu2002performance , Zhang’s PCPzhang2018resource and OnDocLiuyan2019online . HEFT is a classic workflow scheduling algorithm for the heterogeneous computing. It has been widely used for task graph scheduling in different distributed systems. Zhang’s PCP is a task graph scheduling algorithm for the MEC system, which utilizes the partial critical path to reduce the transfer time between dependent tasks. OnDoc is an online task graph scheduling algorithm for multiple applicatoins in the MEC system.
@@
Fig. 6 shows the makespan comparison for handling various applications with $\lambda=\{5,7,9\}$, respectively. We can observe that SATA-DRL achieve the lowest makespan among the four algorithms. Particularly, the makespans generated by HEFT and Zhang’s PCP increase along with the increase of application arrival rates but SATA-DRL and OnDoc are hardly affected by arrival rates. This is because that HEFT and Zhang’s PCP only make scheduling decision for the task among a single application every time without considering the arrvial of multiple applications. Therefore, the application task newly sent from other MUs may still be scheduled to the busy ECDs, resulting in backlogs in the computing queues of ECDs. In addition, the gap between SATA-DRL and OnDoc is relatively small but the makespan achieved by SATA-DRL is less than OnDoc. The reason is that SATA-DRL can learn from the variation of environments and adaptively make more appropriate scheduling decisions for task graph scheduling. Furthermore, we can observe from Fig. 7 that the deadline violation rate of SATA-DRL is much lower than the other algorithms. This further illustrates that SATA-DRL can fully orchestrate the computing resources of the system to make scheduling decisions with careful consideration of the variation of environments.
@@ 7 Conclusions
This paper has investigated the problem of the task graph offloading in MEC, where the computation capabilities of edge computing devices are time-varying. To adapt to environmental changes, we have modeled the task graph scheduling for computation offloading as a Markov Decision Process (MDP). According to the characterization of the environment, we have formulated it as the state space and abstracted the task scheduling decisions into the action space. Moreover, we have defined the reward with respect to MDP as the benefit for the agent. Built upon the task graph scheduling mechanism, we have designed the SATA-DRL algorithm to learn the task scheduling strategy from the interaction with the environment, improving user experience. Extensive experiment results have also validated the superiority of SATA-DRL, by comparing it with existing algorithms, in terms of reducing average makespan and deadline violation.

### v4

@@ 7.3 Evaluation for Reinforcement Learning
In this section, we use cumulative rewards to evaluate the performance of our reinforcement learning. Considering that MUs can randomly send offloading requests of applications, we evaluate cumulative rewards of the reinforcement learning with different arrival rates, i.e., $\lambda=\{5,7,9\}$, when it handles 10 applications. In the experiments, we let 10 applications per episode reach the simulator with different arrival rates, and then launch Algorithm 3 to train the agent and observe the reward values. From Figs. 5, we can observe that the total rewards obtained by the agent are not large at the beginning of the learning process. The total rewards increase as we augment the number of episodes for training. Moreover, the total rewards show relatively stable trends when agent is trained for more than 600 episodes, which is shown in Fig. 5(c). This can be interpreted that the agent can make a relatively optimized decision in every scheduling for the task. As a result, the cumulative rewards can converge. The above results demonstrate the convergence performance of SATA-DRL.
@@ 7.4 Evaluation for Optimization Objective
STATA-DRL is a task graph offloading strategy for multiple applications in the MEC system to minimize the average makespan of applications. There is no fine-grained task graph offloading that utilize DRL to learn the time-varying computing capabilities of ECDs. To evaluate the optimization performance of SATA-DRL, we compare it with three competitors based on heuristic or approximation methods, i.e., Zhang’s PCPzhang2018resource , OnDocLiuyan2019online , and DTO-CED zhang2023dependent , which are also to minimize the makespan of task graph applications MEC system. Moreover, we compare SATA-DRL with a new reinforcement algorithm, named Dueling DQN, to advantages of SATA-DRL. In the experiments that evaluate Dueling DQN, we still utilize the MDP model presented in Section 5.2, that is, the state space, the action space, and the reward are the same as SATA-DRL. The Dueling DQN receives the input of the state space and reward and then outputs the action.
@@
Fig. 6 shows the makespan comparison for handling various applications with $\lambda=\{5,7,9\}$, respectively. We can observe that SATA-DRL achieves the lowest makespan among the five algorithms. Particularly, the makespans generated by Zhang’s PCP increase along with the increase of application arrival rates but the other algorithms are hardly affected by arrival rates. This is because that Zhang’s PCP only make scheduling decision for the task among a single application every time without considering the arrvial of multiple applications. Therefore, the application tasks newly sent from other MUs may still be scheduled to the busy ECDs, resulting in backlogs in the computing queues of ECDs. In addition, the gap among SATA-DRL, DTO-CEO, and OnDoc is relatively small but the makespan achieved by SATA-DRL is less than DTO-CEO and OnDoc. The reason is that SATA-DRL can learn from the variation of environments and adaptively make more appropriate scheduling decisions for task graph scheduling. Furthermore, we can observe from Fig. 7 that the deadline violation rate of SATA-DRL is much lower than the other algorithms. This further illustrates that SATA-DRL can fully orchestrate the computing resources of the system to make scheduling decisions with careful consideration of the variation of environments.
From Figs. 6 and 7, we can also observe that SATA-DRL is superior to Dueling DQN in terms of reducing average makespan and deadline violation. The comparison with the other reinforcement algorithm confirms that the classic DQN method achieves better performance of task graph offloading compared to Dueling DQN.
@@ 8 Conclusions
This paper has investigated the problem of the task graph offloading in MEC, where the computation capabilities of edge computing devices are time-varying. To adapt to environmental changes, we have modeled the task graph scheduling for computation offloading as a MDP. According to the characterization of the environment, we have formulated it as the state space and abstracted the task scheduling decisions into the action space. Moreover, we have defined the reward with respect to MDP as the benefit for the agent. Built upon the task graph scheduling mechanism, we have designed the SATA-DRL algorithm to learn the task scheduling strategy from the interaction with the environment, improving user experience. Extensive experiment results have also validated the superiority of SATA-DRL, by comparing it with existing algorithms, in terms of reducing average makespan and deadline violation.